RSS Feed

Nodes n20-b17-lt1 and s8-a3-lt1 maintenance

Start on 2017-01-09 23:00:00 Status Completed Maintenance Finish on 2017-01-10 00:00:00

Please take a note, that on 2017-01-09 from 23:00 till 00:00 (GMT +2) we will perform maintenance in nodes n20-b17-lt1 and s8-a3-lt1. Due to it all servers in these nodes will be rebooted, expect downtime no more than 20 minutes.

Related servers / services / nodes

* n20-b17-lt
* s8-a3-lt1

KVM servers migration to the new hardware

Start on 2017-01-05 09:00:00 Status Finished Finish on 2017-01-05 12:00:00

Please be advised, that on 2017-01-05 from 09:00 till 2017-01-05 12:00 GMT+2 we will migrate KVM servers to new node hardware.

Due to operation complexity each KVM server on the node can be inaccessible up to 15 minutes.

Related servers / services / nodes

* k23-b15-lt1
* k24-b15-lt1

Node s2-a8-lt1 malfunction

Start on 2017-01-04 03:31:00 Status Resolved Fault / Issue

Please be advised that we had issues with node "s2-a8-lt1". Due to this node was replaced by new one. At this moment all servers are starting.

Related servers / services / nodes

* s2-a8-lt1